Calculate the standard heat of vaporization for tin( IV)chloride, SnCl4.

change heat of fusion=-511.3kj/mol for SnCl4(l) and -471.5 kj for SnCl4 (g).


1
Expert's answer
2022-02-07T17:10:04-0500

Equation 1: Sn(s)+2Cl2(s)→SnCl4(l). ∆H°=-511.3 KJ/mol.....

Equation 2: Sn(s)+2Cl2(g)→SnCl4(l). ∆H=-471.5 KJ/mol....

Equation 2- equation 1

SnCl4(l)-SnCl4(g)

∆H=-471.5-(-511.3)

Ans=+39.8 KJ/mol